What is the pH to the correct number of significant figures for solutions with the following concentrations of H3O+ or OH−?

[OH−] = 5.0×10−3 M pH=

[H3O+] = 10−8 M pH=

[H3O+] = 4.0 M . pH=
Answer #1

[OH−] = 5.0×10−3 M

pH= 14 - Poh

Poh = -log[OH-] = -log(5.0*10^-3) = 2.3

Ph = 14 - 2.3 = 11.7

Ph = 12

,...........................................................................

[H3O+] = 10−8 M

pH= -log[H3O+]

Ph = -log[10^-8]

Ph = 8

[H3O+] = 4.0 M .

pH= -log[H3O+]

Ph = -log[4]